# A Survey on Multi-Generative Agent System: Recent Advances and New Frontiers

### v1

@@ 5.2 Resources of MGAS for evaluations
Table 3: Codes and Benchmarks in MGAS for evaluation studies. “No Code” or “No Benchmark” means the code or benchmark is unavailable.
@@
Training on generative agents SFT on MGAS Li et al. (2024c) Code Link MT-Bench, AlpacaEval
MARL on MGAS Xu et al. (2023c) No Code No dataset or benchmark
@@ 6 Challenges and Future Directions
While previous work on MGAS has obtained many remarkable successes, this field is still at its initial stage, and there are several significant challenges that need to be addressed in its development. In the following, we outline several key challenges along with potential future directions.
@@ 6.1 Challenges posed by generative agents
Generative agents are an integral part of MGAS. However, the generative agents have some shortcomings due to the inherent characteristics of the base model LLMs, which will be carefully discussed below.
@@ 6.2 Challenges posed by interactions
Due to the complexity, autoregressive, and other characteristics of MGAS, there are many problems in the practical application of the system. Two main problems, Efficiency explosion, and Accumulative Effect, are listed in the following.
Efficiency explosion. Due to their autoregressive architecture, LLMs typically have slow inference speeds. However, generative agents need to query LLMs for each action multiple times, such as extracting information from memory, making plans before taking actions, and so on. When the MGAS scales up, this problem will be magnified, especially for generative agents that have a large action space. SoMoSiMu-Bench Mou et al. (2024) replaces the edge generative agents with rule-based agents, alleviating this problem. However, for MGAS with complex action apace in generative agents, this problem remains unsolved.
Accumulative Effect. Since each round of MGAS is based on the results of the previous round, and MGAS has no means of error correction, each error or error will have a great impact on the subsequent results. Researchers have used a rule-based model for intermediate error correctionChen et al. (2024c), but there is still a lot of room for improvement. IOA Chen et al. (2024f) proposed an Internet-like communication architecture, which made mGAS more scalable and enhanced the adaptability to dynamic tasks.
Future directions. Industry academia has been making efforts to reduce the communication cost of MGAS, such as alignment-based method OPTIMA Chen et al. (2024g) and Industrialized parallel message method AgentScope Gao et al. (2024), but it is still in the basic stage and has a large research space.
6.3 Challenges of Evaluating for MGAS
Lack of Objective metrics for group behavior. As shown in Section 4.3, due to the diversity, complexity, and unpredictability of multi-agent environments, it is difficult to obtain sufficiently detailed, specific, and direct system evaluation indicators from current work at the population level. At present, researchers mainly compare the distribution of the system and real environments to evaluate MGAS, which lacks details for the MGAS running process.
Automated evaluation and benchmark. Different MGAS of the same kind can not be compared because of the lack of a benchmark for MGAS. Further, There is a lack of a common benchmark framework for both individual and total-based evaluation, that can be used to evaluate most MGAS.
Future directions. Studying large-scale MGAS will be a new research hotspot, from which researchers will evaluate and discover new scale effects. In the meantime, common test benchmarks and evaluation methods will also emerge in future research.
@@ 7 Conclusion
In this survey, we systematically summarize existing research in the field of Multi-generative agent system (MGAS). We present and review these studies from three application aspects including task-solving, simulation, and evaluation of the MGAS. For each of these aspects, we provide a detailed taxonomy to draw connections among the existing research, summarizing the major techniques and their development histories. In addition to reviewing the previous work, we also propose several challenges in this field, which are expected to guide potential future directions.
@@
We have made our best effort, but some limitations may still exist. Due to page limitations, we can only provide a brief summary of each method without exhaustive technical details. On the other hand, we primarily collect studies from *ACL, NeurIPS, ICLR, AAAI, and arXiv, and there is a chance that we may have missed some important work published in other venues. In application, we primarily list representative MGAS resources with open code in Table 1, Table 2, and Table 3. More complete papers can be found in https://github.com/bianhua-12/Multi-generative_Agent_System_survey. We recognize the timeliness of our work, and we will stay abreast of discussions within the research community, updating opinions and supplementing overlooked work in the future.

### v2

@@ 5.2 Resources of LLM-MAS for evaluations
Table 3: Codes and Benchmarks in LLM-MAS for evaluation studies. “No Code” or “No Benchmark” means the code or benchmark is unavailable.
@@
Training on generative agents SFT on LLM-MAS Li et al. (2024c) Code Link MT-Bench, AlpacaEval
MARL on LLM-MAS Xu et al. (2023c) No Code No dataset or benchmark
@@ 6 Challenges and Future Directions
While previous work on LLM-MAS has obtained many remarkable successes, this field is still at its initial stage, and there are several significant challenges that need to be addressed in its development. In the following, we outline several key challenges along with potential future directions.
@@ 6.1 Challenges posed by generative agents
Generative agents are an integral part of LLM-MAS. However, the generative agents have some shortcomings due to the inherent characteristics of the base model LLMs, which will be carefully discussed below.
@@ 6.2 Challenges posed by interactions
Due to the complexity, autoregressive, and other characteristics of LLM-MAS, there are many problems in the practical application of the system. Two main problems, Efficiency explosion, and Accumulative Effect, are listed in the following.
Efficiency explosion. Due to their autoregressive architecture, LLMs typically have slow inference speeds. However, generative agents need to query LLMs for each action multiple times, such as extracting information from memory, making plans before taking actions, and so on. When the LLM-MAS scales up, this problem will be magnified, especially for generative agents that have a large action space. SoMoSiMu-Bench Mou et al. (2024) replaces the edge generative agents with rule-based agents, alleviating this problem. However, for LLM-MAS with a complex action space in generative agents, this problem remains unsolved.
Accumulative Effect. Since each round of LLM-MAS is based on the results of the previous round, and LLM-MAS have a great impact on the subsequent results. Researchers have used a rule-based model for intermediate error correction Chen et al. (2024c), but there is still a lot of room for improvement. IOA Chen et al. (2024f) proposed an Internet-like communication architecture, which made LLM-MAS more scalable and enhanced the adaptability to dynamic tasks.
Future directions. Industry academia has been making efforts to reduce the communication cost of LLM-MAS, such as alignment-based method OPTIMA Chen et al. (2024g) and Industrialized parallel message method AgentScope Gao et al. (2024), but it is still in the basic stage and has a large research space.
6.3 Challenges of Evaluating for LLM-MAS
Lack of Objective metrics for group behavior. As shown in Section 4.3, due to the diversity, complexity, and unpredictability of multi-agent environments, it is difficult to obtain sufficiently detailed, specific, and direct system evaluation indicators from current work at the population level. At present, researchers mainly compare the distribution of the system and real environments to evaluate LLM-MAS, which lacks details for the LLM-MAS running process.
Automated evaluation and benchmark. Different LLM-MAS of the same kind cannot be compared because of the lack of a benchmark for LLM-MAS. Further, there is a lack of a common benchmark framework for both individual and total-based evaluation, that can be used to evaluate most LLM-MAS.
Future directions. Studying large-scale LLM-MAS will be a new research hotspot, from which researchers will evaluate and discover new scale effects. In the meantime, common test benchmarks and evaluation methods will also emerge in future research.
@@ 7 Conclusion
In this survey, we systematically summarize existing research in the LLM-based Multi-Agent Systems (LLM-MAS) field. We present and review these studies from three application aspects: task-solving, simulation, and evaluation of the LLM-MAS. We provide a detailed taxonomy to draw connections among the existing research, summarizing the major techniques and their development histories for each of these aspects. In addition to reviewing the previous work, we also propose several challenges in this field, which are expected to guide potential future directions.
@@
We have made our best effort, but some limitations may still exist. Due to page limitations, we can only provide a brief summary of each method without exhaustive technical details. On the other hand, we primarily collect studies from *ACL, NeurIPS, ICLR, AAAI, and arXiv, and there is a chance that we may have missed some important work published in other venues. In application, we primarily list representative LLM-MAS resources with open code in Table 1, Table 2, and Table 3. More complete papers can be found in https://github.com/bianhua-12/Multi-generative_Agent_System_survey. We recognize the timeliness of our work, and we will stay abreast of discussions within the research community, updating opinions and supplementing overlooked work in the future.